Circular No. 6557

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997Y IN NGC 4675 A. V. Filippenko, D. C. Leonard, A. M. Gilbert, and W. C. G. Ho,

University of California at Berkeley, report that a CCD spectrum (range 430–700 nm) obtained on Feb. 9 with the Shane 3-m reflector at Lick Observatory shows that SN 1997Y (cf. IAUC 6556) is of type Ia, close to maximum brightness. The spectrum resembles that of the prototypical type-Ia supernova 1981B. There is prominent Na I D interstellar absorption (equivalent width roughly 0.13 nm), indicating that the object is significantly reddened.

P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, write: “A spectrum of SN 1997Y was obtained by J. Mader on Feb. 9.4 UT with the 1.5-m Tillinghast telescope. The spectrum shows that the object is a type-Ia supernova at maximum brightness. Narrow emission lines from the galaxy give a redshift of z = 0.0162. An absorption line at 598.71 nm is identified as interstellar Na I in the host galaxy and has an equivalent width of 0.16 nm. CCD images obtained by D. Kim with the Whipple Observatory 1.2-m telescope on Feb. 9.5 show SN 1997Y to be 3′′.8 west and 1′′.7 north of the galaxy’s nucleus, at V = 15.3 ± 0.1.”

SUPERNOVAE 1996cd AND 1997T Filippenko et al., observing as above, also report that SN 1996cd (cf.

IAUC 6531) is indeed a supernova, probably of type Ib or Ic. No broad features are seen at the expected positions of Hα or Si II 635.5-nm, but there is a strong absorption trough at the position of He I 587.6-nm (or Na I D). The redshift of the object is 0.048, determined from absorption lines in the nucleus of the host galaxy and from narrow emission lines at the position of the supernova. SN 1997T (cf. IAUC 6543) is of type Ia, roughly one month past maximum brightness; the redshift of the host galaxy is 0.042, determined from narrow emission lines in the nucleus.

SUPERNOVAE 1997W AND 1996bw IN NGC 664 C. Hergenrother, Lunar and Planetary Laboratory, reports the follow-

ing position for SN 1997W from the discovery image (cf. IAUC 6551): α = 1^{h}43^{m}44^{s}.48, δ = +4^{o}13′17′′.1 (equinox 2000.0). The position end figures for SN 1996bw are 44^{s}.51, 19′′.0 (cf. IAUC 6516).

1997 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green